OpenStudy (anonymous):

simplify (9x^7/15x^5)(5x^2/45x^3)

OpenStudy (anonymous):

\[(\frac{ 9x^7 }{ 15x^5 })(\frac{ 5x^2 }{ 45x^3 })\]

OpenStudy (anonymous):

step by step please

OpenStudy (anonymous):

you can cancel an x^2 out of all of the terms leaving you with ((9x^5)/(15x^3))(5)/(45x)) multiply the numerators and the denominators (45x^5)/(645x^4) reduce; 45/645= 1/15 and x^5/x^4=x^1 x/15

OpenStudy (anonymous):

\[\Large (\frac{9x^7}{15x^5})(\frac{5x^2}{45x^3})\]\[\Large (\frac{3x^7}{5x^5})(\frac{x^2}{9x^3})\]\[\Large \frac{(3x^7)(x^2)}{(5x^5)(9x^3)}\]\[\Large \frac{3x^9}{45x^8}\]\[\Large \frac{x^9}{15x^8}\]\[\Large \frac{x}{15}\]

OpenStudy (anonymous):

Oh sorry I didn't see that there was already an answer